# Compositional Monte Carlo Tree Diffusion for Extendable Planning

### v1

@@ 1 Introduction
Diffusion models have emerged as a powerful framework for trajectory planning, particularly excelling in long-horizon tasks due to their ability to learn plans in a holistic manner [15, 2, 23]. This stands in contrast to conventional autoregressive approaches [11, 12, 3], which construct trajectories step-by-step. By generating trajectories as coherent wholes, diffusion planners can effectively address key challenges in long-term planning, such as sparse rewards and the accumulation of transition errors. However, these models often struggle to generate complex long-horizon plans rarely encountered in the training distribution.
To address this limitation, a growing body of work has explored inference-time scaling techniques, particularly by incorporating them into the diffusion denoising process [25, 27, 32, 33]. Among these approaches, Monte Carlo Tree Diffusion (MCTD) [32] represents a novel perspective by combining tree search with denoising. This integration enables MCTD to achieve state-of-the-art performance across a wide range of planning tasks.
MCTD combines causally scheduled denoising with Monte Carlo Tree Search (MCTS) [7], establishing a structured framework for deliberative planning that has shown strong performance across diverse planning domains. By decomposing trajectory generation into a subplan-level tree process, which decomposes a full plan into multiple subplans and systematically searches for the best combination of these subplans performing exploration and exploitation to escape local optima. It has demonstrated notable success in complex, long-horizon planning tasks [28] where conventional diffusion planners often struggle, representing a substantial advancement in generative planning methods.
Despite its strong performance, MCTD inherently struggles to generate plans that exceed the trajectory lengths observed during training. While existing periodic replanning strategies appear to address this length limitation, they suffer from a fundamental myopic decision-making problem that severely undermines their effectiveness [32]. Specifically, the replanning approaches make planning decisions based solely on local information within each limited planning horizon. It can lead to dead-end states or suboptimal paths due to the lack of consideration about the global path structure or long-term consequences. This fundamental limitation raises a key research question: How can we enable MCTD to perform globally-aware planning that substantially exceeds training trajectory lengths while avoiding the myopic pitfalls of traditional replanning approaches?
@@ 2.1 Diffusion models for planning
Diffusion models have been successfully applied to planning by reformulating it as a generative modeling problem [15, 2, 34, 1, 23], where trajectories are represented as sequential data comprising state-action pairs. Formally, a trajectory is denoted as $\mathbf{x}=[s_{0};a_{0},s_{1};a_{1},\dots,s_{T};a_{T}]$, where $T$ is the planning horizon and $(s_{t},a_{t})$ represents the state-action pair at time step $t$. During inference, the trajectory is iteratively refined from an initial random noise through a sequence of denoising steps, progressively converging toward a coherent plan.
This approach differs fundamentally from traditional planning algorithms like Probabilistic Road Maps (PRM) [16] and Rapidly-exploring Random Trees (RRT) [20]. Whereas traditional methods are online algorithms that actively explore the state space by querying an environment simulator or interacting with the physical world, diffusion planners operate in a fully offline setting. They learn to generate the trajectories from a static dataset of prior experiences, requiring no online environment access during the planning phase.
To ensure that generated trajectories align with specified task objectives, both classifier-free [14] and classifier-guided sampling [8] techniques have been explored in the context of diffusion planning [15, 2, 1]. In this work, we adopt the classifier-guided approach [15], which incorporates a guidance function $\mathcal{J}_{\phi}(\mathbf{x})$ to steer the generative process toward high-reward trajectories. The modified sampling distribution is defined as:
@@
Additionally, we adopt semi-autoregressive generation with causal noise schedule [2], which selectively denoises uncertain tokens (typically future timesteps) while preserving causal dependencies. This approach maintains temporal consistency and significantly improves long-horizon plan quality compared to standard diffusion generation.
@@ 2.2 Monte Carlo Tree Diffusion
Monte Carlo Tree Diffusion (MCTD) [32] reformulates trajectory planning as a tree search problem, structured around three key components.
Denoising as tree rollout. Unlike conventional MCTS approaches that perform rollouts at the individual state level, MCTD operates at the level of subplans—partitioned segments of the full trajectory. A trajectory is represented as $\mathbf{x}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{S}]$, where $\mathbf{x}$ denotes the complete trajectory and each $\mathbf{x}_{s}$ is a constituent subplan. MCTD leverages a semi-autoregressive generation strategy [2], resulting in the following factorization: $\smash{p(\mathbf{x})\approx\prod_{s=1}p(\mathbf{x}_{s}|\mathbf{x}_{1:s-1})}$. This formulation preserves the global coherence characteristic of diffusion models while enabling intermediate evaluations akin to MCTS rollouts, effectively reducing the depth of the search tree.
@@
Jumpy denoising as fast simulation. To enable efficient simulation, MCTD leverages fast denoising techniques such as Denoising Diffusion Implicit Models (DDIM) [30] to complete the remaining noisy subplans using fewer denoising steps. Given a partially generated trajectory $\mathbf{x}_{1:s}$, the remaining segments $\tilde{\mathbf{x}}_{s+1:S}$ are rapidly denoised as: $\tilde{\mathbf{x}}_{s+1:S}\sim p(\mathbf{x}_{s+1:S}|\mathbf{x}_{1:s},g)$, where $g$ denotes the guidance level. This produces a complete trajectory $\tilde{\mathbf{x}}$ that can be used for evaluation, enabling fast and approximate rollout within the tree.
@@
(1) Selection. The tree is traversed from the root using a node selection policy such as UCT [19]. Unlike standard MCTS, each node in MCTD corresponds to an extended subplan rather than a single state, reducing tree depth and promoting higher-level abstraction. (2) Expansion. When a leaf node is reached, a meta-action—the guidance level gg—is selected. A new subplan $\mathbf{x}_{s}$ is then sampled conditioned on $g$. (3) Simulation. After expansion, the remaining subplans are rapidly completed using accelerated denoising techniques such as DDIM [30]. Although approximate, this simulation step offers significant computational efficiency while retaining sufficient fidelity for reward-based evaluation using the reward function, $r(\mathbf{x})$. (4) Backpropagation. The resulting statistics, including visitation counts and trajectory rewards, are propagated upward through the tree to update value estimates and guide future search.
@@ 3.1 Online Composer
To implement plan-level tree search, OC performs stitching at each tree expansion by connecting the terminal state of the parent node’s plan to the starting state of the newly generated plan. This enables each expansion to generate longer plans as the length of the plans generated from the diffusion planner [15, 2, 32] while systematically searching for global optima through tree exploration. The resulting trajectory is a stitched sequence of plans $\mathbf{x}=[\mathbf{x}^{1},\dots,\mathbf{x}^{M}]$, generated autoregressively:
@@
In MCTD [32], the guidance level—a parameter controlling the influence of a guidance function on generation—is used as a meta-action for branching the search tree. Our approach generalizes this concept to the plan level by introducing a guidance set as its meta-action. This set, comprising multiple guidance levels, enables inference-time planners (e.g., Best-of-$N$ or MCTD) to dynamically select the most appropriate level at each step. This mechanism empowers the planner to generate more effective and diverse local plans by adaptively modulating the guidance function’s influence during the tree search. We empirically demonstrate the efficacy of using guidance sets in Appendix B.1. By enabling more sophisticated plan generation at each compositional step, this approach significantly enhances the quality of the final stitched solutions.
@@
To accelerate planning, OC employs fast replanning inspired by jumpy denoising [30] from MCTD. Once tree expansion reaches the $m$-th plan, the remaining trajectory is rapidly completed through a replanning process with the jumpy denoising that skips the denoising step every $C$ steps:
@@
While OC effectively extends planning beyond training trajectory lengths, it faces scalability challenges in extremely large spaces due to exponential search growth inherent in sequential tree search [7, 32]. These limitations are empirically demonstrated in our maze experiments (Section 5.2). To address these challenges, we introduce Distributed Composer and Preplan Composer in the following sections.
@@ 3.2 Distributed Composer
After parallel tree expansion and connection, DC constructs a connectivity graph $G$ where nodes are the starting positions $\mathcal{S}$ and edges represent feasible plans between them. DC then applies classical shortest-path algorithms (Dijkstra [9] or A* [13]) to find optimal plan combinations. Search terminates when the synthesized plan satisfies task constraints (e.g., maximum episode length), ensuring computational efficiency by avoiding unnecessary exploration once a viable solution is found.
@@ 3.3 Preplan Composer
When presented with a goal-conditioned planning task $(s_{\text{start}},s_{\text{goal}})$, PC efficiently leverages the prebuilt plan graph for solution composition. The algorithm employs Online Composer to generate only short connecting plans: from $s_{\text{start}}$ to waypoints and from waypoints to $s_{\text{goal}}$. Since long-distance connections between waypoints are already planned in the preprocessing phase, OC focuses solely on these local connections. Using the augmented graph, PC applies shortest-path algorithms [9, 13] to identify the optimal waypoint sequence forming a complete solution. This approach dramatically reduces search complexity compared to DC’s extensive online search, as most potential paths are pre-encoded in the graph.
@@ 4 Related work
Diffusion models [29] excel at long-horizon planning, particularly for sparse reward settings, by learning to generate plans holistically [15, 2, 1, 34, 23]. Research advances include hierarchical planning approaches for improved efficiency and long-term reasoning [5, 6, 10], hybrid planners integrating value learning policies [4], and diffusion planning with causal noise schedule for generating causally consistent plans [2].
@@
Recent works have explored inference-time scaling to enhance diffusion planner reasoning capabilities [32, 33]. Yoon et al. [32] integrates semi-autoregressive denoising [2] with Monte Carlo Tree Search (MCTS) [7], while Zhang et al. [33] applies MCTS to diffusion models using learnable energy functions for value estimation.
@@
Several approaches extend planning capabilities beyond training trajectory lengths [21, 6, 17, 24], primarily by modifying the training data or objectives. For instance, some methods create longer training sequences by stitching trajectories from the dataset before training the diffusion planner [21, 6]. Others introduce specialized training objectives that encourage the model to learn compositional structures from sub-trajectories [24], or employ dedicated value functions for long-term estimation while generating short-horizon plans [17].
@@ 5 Experiments
To demonstrate the effectiveness of C-MCTD in compositional long-horizon planning, we conduct a comprehensive evaluation on tasks from the Offline Goal-conditioned RL benchmark (OGBench) [28], following MCTD’s experimental setup [32]. Our evaluation spans point and ant maze navigation with extended horizons, multi-cube robot arm manipulation, and partially observable visual maze tasks. All evaluated methods are diffusion planners, which, as discussed in Section 2.1, are trained on offline trajectory datasets and do not require environment interaction during training or planning. We report mean success rate (%) and planning time (seconds), averaged over 50 runs (5 tasks × 10 seeds), with complete configuration details in Appendix A.
@@ 5.1 Baselines
We compare C-MCTD against diverse baselines spanning stitching methods and diffusion planners. For stitching approaches, we evaluate Replan (fixed-interval plan regeneration without plan-level search) and DatasetStitch [21, 6] (training on concatenated trajectories). For diffusion planners, we include Diffuser [15] (foundational diffusion planning), Diffusion Forcing [2] (causal noise scheduling for long-horizon control), SSD [17] (stitching diffusion planner via learned value function guidance), CompDiffuser [24] (stitching diffusion planner with sub-trajectory dependency modeling), and MCTD [32] (direct comparison for plan-level vs. in-plan search). Complete details are provided in Appendix A.3.
@@ 5.2 Long-horizon maze navigation: scaling beyond trained trajectory length
We evaluate the effectiveness of C-MCTD on long-horizon tasks using PointMaze and AntMaze environments from OGBench, which require planning trajectories up to 1000 steps despite training on significantly shorter sequences (100 steps). Following prior work [15, 32], PointMaze uses a heuristic controller while AntMaze employs a learned value-based policy [31].
@@ 5.3 Multi-object manipulation: compositional planning with plan caching
We evaluate C-MCTD on multi-cube manipulation tasks from OGBench [28], which require precise action sequencing for stacking cubes into predefined configurations. These tasks demand compositional reasoning to determine optimal cube movement sequences, making them ideal for testing plan-level search capabilities. Our approach combines diffusion-based high-level planning with low-level value-guided control and object-wise guidance as described in prior work [4, 32].
@@
Baseline performance degrades with task complexity. Table 3 shows that while MCTD-Replan outperforms Diffuser-based approaches due to object-wise guidance capabilities [32], its effectiveness diminishes substantially as complexity increases—dropping from 100% (single cube) to 24% (quadruple cube). This performance degradation highlights the limitations of the baseline’s search strategy in complex sequential manipulation tasks.

### v2

@@ 1 Introduction
Diffusion models have emerged as a powerful framework for trajectory planning, particularly excelling in long-horizon tasks due to their ability to learn plans in a holistic manner [diffuser, diffusion_forcing, diffusion_planner_invest]. This stands in contrast to conventional autoregressive approaches [worldmodel, dreamer, transdreamer], which construct trajectories step-by-step. By generating trajectories as coherent wholes, diffusion planners can effectively address key challenges in long-term planning, such as sparse rewards and the accumulation of transition errors. However, these models often struggle to generate complex long-horizon plans rarely encountered in the training distribution.
To address this limitation, a growing body of work has explored inference-time scaling techniques, particularly by incorporating them into the diffusion denoising process [sop, dlbs, mctd, t_scend]. Among these approaches, Monte Carlo Tree Diffusion (MCTD) [mctd] represents a novel perspective by combining tree search with denoising. This integration enables MCTD to achieve state-of-the-art performance across a wide range of planning tasks.
MCTD combines causally scheduled denoising with Monte Carlo Tree Search (MCTS) [mcts], establishing a structured framework for deliberative planning that has shown strong performance across diverse planning domains. By decomposing trajectory generation into a subplan-level tree process, which decomposes a full plan into multiple subplans and systematically searches for the best combination of these subplans performing exploration and exploitation to escape local optima. It has demonstrated notable success in complex, long-horizon planning tasks [ogbench] where conventional diffusion planners often struggle, representing a substantial advancement in generative planning methods.
Despite its strong performance, MCTD inherently struggles to generate plans that exceed the trajectory lengths observed during training. While existing periodic replanning strategies appear to address this length limitation, they suffer from a fundamental myopic decision-making problem that severely undermines their effectiveness [mctd]. Specifically, the replanning approaches make planning decisions based solely on local information within each limited planning horizon. It can lead to dead-end states or suboptimal paths due to the lack of consideration about the global path structure or long-term consequences. This fundamental limitation raises a key research question: How can we enable MCTD to perform globally-aware planning that substantially exceeds training trajectory lengths while avoiding the myopic pitfalls of traditional replanning approaches?
@@ 2.1 Diffusion models for planning
Diffusion models have been successfully applied to planning by reformulating it as a generative modeling problem [diffuser, diffusion_forcing, d-mpc, decisiondiffuer, diffusion_planner_invest], where trajectories are represented as sequential data comprising state-action pairs. Formally, a trajectory is denoted as $\mathbf{x}=[s_{0};a_{0},s_{1};a_{1},\dots,s_{T};a_{T}]$, where $T$ is the planning horizon and $(s_{t},a_{t})$ represents the state-action pair at time step $t$. During inference, the trajectory is iteratively refined from an initial random noise through a sequence of denoising steps, progressively converging toward a coherent plan.
This approach differs fundamentally from traditional planning algorithms like Probabilistic Road Maps (PRM) [prm] and Rapidly-exploring Random Trees (RRT) [rrt]. Whereas traditional methods are online algorithms that actively explore the state space by querying an environment simulator or interacting with the physical world, diffusion planners operate in a fully offline setting. They learn to generate the trajectories from a static dataset of prior experiences, requiring no online environment access during the planning phase.
To ensure that generated trajectories align with specified task objectives, both classifier-free [classifier_free] and classifier-guided sampling [classifier_guided] techniques have been explored in the context of diffusion planning [diffuser, diffusion_forcing, decisiondiffuer]. In this work, we adopt the classifier-guided approach [diffuser], which incorporates a guidance function $\mathcal{J}_{\phi}(\mathbf{x})$ to steer the generative process toward high-reward trajectories. The modified sampling distribution is defined as:
@@
Additionally, we adopt semi-autoregressive generation with causal noise schedule [diffusion_forcing], which selectively denoises uncertain tokens (typically future timesteps) while preserving causal dependencies. This approach maintains temporal consistency and significantly improves long-horizon plan quality compared to standard diffusion generation.
@@ 2.2 Monte Carlo Tree Diffusion
Monte Carlo Tree Diffusion (MCTD) [mctd] reformulates trajectory planning as a tree search problem, structured around three key components.
Denoising as tree rollout. Unlike conventional MCTS approaches that perform rollouts at the individual state level, MCTD operates at the level of subplans—partitioned segments of the full trajectory. A trajectory is represented as $\mathbf{x}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{S}]$, where $\mathbf{x}$ denotes the complete trajectory and each $\mathbf{x}_{s}$ is a constituent subplan. MCTD leverages a semi-autoregressive generation strategy [diffusion_forcing], resulting in the following factorization: $\smash{p(\mathbf{x})\approx\prod_{s=1}p(\mathbf{x}_{s}|\mathbf{x}_{1:s-1})}$. This formulation preserves the global coherence characteristic of diffusion models while enabling intermediate evaluations akin to MCTS rollouts, effectively reducing the depth of the search tree.
@@
Jumpy denoising as fast simulation. To enable efficient simulation, MCTD leverages fast denoising techniques such as Denoising Diffusion Implicit Models (DDIM) [ddim] to complete the remaining noisy subplans using fewer denoising steps. Given a partially generated trajectory $\mathbf{x}_{1:s}$, the remaining segments $\tilde{\mathbf{x}}_{s+1:S}$ are rapidly denoised as: $\tilde{\mathbf{x}}_{s+1:S}\sim p(\mathbf{x}_{s+1:S}|\mathbf{x}_{1:s},g)$, where $g$ denotes the guidance level. This produces a complete trajectory $\tilde{\mathbf{x}}$ that can be used for evaluation, enabling fast and approximate rollout within the tree.
@@
(1) Selection. The tree is traversed from the root using a node selection policy such as UCT [uct]. Unlike standard MCTS, each node in MCTD corresponds to an extended subplan rather than a single state, reducing tree depth and promoting higher-level abstraction. (2) Expansion. When a leaf node is reached, a meta-action—the guidance level gg—is selected. A new subplan $\mathbf{x}_{s}$ is then sampled conditioned on $g$. (3) Simulation. After expansion, the remaining subplans are rapidly completed using accelerated denoising techniques such as DDIM [ddim]. Although approximate, this simulation step offers significant computational efficiency while retaining sufficient fidelity for reward-based evaluation using the reward function, $r(\mathbf{x})$. (4) Backpropagation. The resulting statistics, including visitation counts and trajectory rewards, are propagated upward through the tree to update value estimates and guide future search.
@@ 3.1 Online Composer
To implement plan-level tree search, OC performs stitching at each tree expansion by connecting the terminal state of the parent node’s plan to the starting state of the newly generated plan. This enables each expansion to generate longer plans as the length of the plans generated from the diffusion planner [diffuser, diffusion_forcing, mctd] while systematically searching for global optima through tree exploration. The resulting trajectory is a stitched sequence of plans $\mathbf{x}=[\mathbf{x}^{1},\dots,\mathbf{x}^{M}]$, generated autoregressively:
@@
In MCTD [mctd], the guidance level—a parameter controlling the influence of a guidance function on generation—is used as a meta-action for branching the search tree. Our approach generalizes this concept to the plan level by introducing a guidance set as its meta-action. This set, comprising multiple guidance levels, enables inference-time planners (e.g., Best-of-$N$ or MCTD) to dynamically select the most appropriate level at each step. This mechanism empowers the planner to generate more effective and diverse local plans by adaptively modulating the guidance function’s influence during the tree search. We empirically demonstrate the efficacy of using guidance sets in Appendix B.1. By enabling more sophisticated plan generation at each compositional step, this approach significantly enhances the quality of the final stitched solutions.
@@
To accelerate planning, OC employs fast replanning inspired by jumpy denoising [ddim] from MCTD. Once tree expansion reaches the $m$-th plan, the remaining trajectory is rapidly completed through a replanning process with the jumpy denoising that skips the denoising step every $C$ steps:
@@
While OC effectively extends planning beyond training trajectory lengths, it faces scalability challenges in extremely large spaces due to exponential search growth inherent in sequential tree search [mcts, mctd]. These limitations are empirically demonstrated in our maze experiments (Section 5.2). To address these challenges, we introduce Distributed Composer and Preplan Composer in the following sections.
@@ 3.2 Distributed Composer
After parallel tree expansion and connection, DC constructs a connectivity graph $G$ where nodes are the starting positions $\mathcal{S}$ and edges represent feasible plans between them. DC then applies classical shortest-path algorithms (Dijkstra [dijkstra] or A* [a_star]) to find optimal plan combinations. Search terminates when the synthesized plan satisfies task constraints (e.g., maximum episode length), ensuring computational efficiency by avoiding unnecessary exploration once a viable solution is found.
@@ 3.3 Preplan Composer
When presented with a goal-conditioned planning task $(s_{\text{start}},s_{\text{goal}})$, PC efficiently leverages the prebuilt plan graph for solution composition. The algorithm employs Online Composer to generate only short connecting plans: from $s_{\text{start}}$ to waypoints and from waypoints to $s_{\text{goal}}$. Since long-distance connections between waypoints are already planned in the preprocessing phase, OC focuses solely on these local connections. Using the augmented graph, PC applies shortest-path algorithms [dijkstra, a_star] to identify the optimal waypoint sequence forming a complete solution. This approach dramatically reduces search complexity compared to DC’s extensive online search, as most potential paths are pre-encoded in the graph.
@@ 4 Related work
Diffusion models [diffusion] excel at long-horizon planning, particularly for sparse reward settings, by learning to generate plans holistically [diffuser, diffusion_forcing, decisiondiffuer, d-mpc, diffusion_planner_invest]. Research advances include hierarchical planning approaches for improved efficiency and long-term reasoning [hd, hmd, diffuserlite], hybrid planners integrating value learning policies [plandq], and diffusion planning with causal noise schedule for generating causally consistent plans [diffusion_forcing].
@@
Recent works have explored inference-time scaling to enhance diffusion planner reasoning capabilities [mctd, t_scend]. mctd integrates semi-autoregressive denoising [diffusion_forcing] with Monte Carlo Tree Search (MCTS) [mcts], while t_scend applies MCTS to diffusion models using learnable energy functions for value estimation.
@@
Several approaches extend planning capabilities beyond training trajectory lengths [diffstitch, hmd, ssd, compdiffuser], primarily by modifying the training data or objectives. For instance, some methods create longer training sequences by stitching trajectories from the dataset before training the diffusion planner [diffstitch, hmd]. Others introduce specialized training objectives that encourage the model to learn compositional structures from sub-trajectories [compdiffuser], or employ dedicated value functions for long-term estimation while generating short-horizon plans [ssd].
@@ 5 Experiments
To demonstrate the effectiveness of C-MCTD in compositional long-horizon planning, we conduct a comprehensive evaluation on tasks from the Offline Goal-conditioned RL benchmark (OGBench) [ogbench], following MCTD’s experimental setup [mctd]. Our evaluation spans point and ant maze navigation with extended horizons, multi-cube robot arm manipulation, and partially observable visual maze tasks. All evaluated methods are diffusion planners, which, as discussed in Section 2.1, are trained on offline trajectory datasets and do not require environment interaction during training or planning. We report mean success rate (%) and planning time (seconds), averaged over 50 runs (5 tasks × 10 seeds), with complete configuration details in Appendix A.
@@ 5.1 Baselines
We compare C-MCTD against diverse baselines spanning stitching methods and diffusion planners. For stitching approaches, we evaluate Replan (fixed-interval plan regeneration without plan-level search) and DatasetStitch [diffstitch, hmd] (training on concatenated trajectories). For diffusion planners, we include Diffuser [diffuser] (foundational diffusion planning), Diffusion Forcing [diffusion_forcing] (causal noise scheduling for long-horizon control), SSD [ssd] (stitching diffusion planner via learned value function guidance), CompDiffuser [compdiffuser] (stitching diffusion planner with sub-trajectory dependency modeling), and MCTD [mctd] (direct comparison for plan-level vs. in-plan search). Complete details are provided in Appendix A.3.
@@ 5.2 Long-horizon maze navigation: scaling beyond trained trajectory length
We evaluate the effectiveness of C-MCTD on long-horizon tasks using PointMaze and AntMaze environments from OGBench, which require planning trajectories up to 1000 steps despite training on significantly shorter sequences (100 steps). Following prior work [diffuser, mctd], PointMaze uses a heuristic controller while AntMaze employs a learned value-based policy [dql].
@@ 5.3 Multi-object manipulation: compositional planning with plan caching
We evaluate C-MCTD on multi-cube manipulation tasks from OGBench [ogbench], which require precise action sequencing for stacking cubes into predefined configurations. These tasks demand compositional reasoning to determine optimal cube movement sequences, making them ideal for testing plan-level search capabilities. Our approach combines diffusion-based high-level planning with low-level value-guided control and object-wise guidance as described in prior work [plandq, mctd].
@@
Baseline performance degrades with task complexity. Table 3 shows that while MCTD-Replan outperforms Diffuser-based approaches due to object-wise guidance capabilities [mctd], its effectiveness diminishes substantially as complexity increases—dropping from 100% (single cube) to 24% (quadruple cube). This performance degradation highlights the limitations of the baseline’s search strategy in complex sequential manipulation tasks.